OpenStudy (anonymous):

Area of the square P is 36 square inches and the perimeter of the rectangle R is 144 inches. If a width of the rectangle is two times a length of the square, which is the length of a length of the rectangle?

OpenStudy (anonymous):

hi there

OpenStudy (anonymous):

first each side of square is 6 ... so rectangle width would be 2*6=12 ... perimeter is 2 * W + 2 * L ... 144 = 2*12 + 2*L .. now its simply a solve for L which I leave for you do battle - enjoy

OpenStudy (anonymous):

Area P=36 in^2 which means: The length of P=sqrt(36)=6 in Perim R=144 in Width R=2*length of P=2*6=12 in Length R=(Perim R-(2*width R))/2=(144-(2*12))/2=60 in

OpenStudy (anonymous):

The wording of the end of your question confuses me but let's see. Since the are is 36, that means that all sides of the square are 6 inches each. If the width of the rectangle is the 2x the length of a side of a square, that's 12. There are two sides so that's 24. 144-24 is 120 and dividing that in 2 would be 60. If this is what you're asking.

OpenStudy (anonymous):

Yes, I got it thanks I didn't include the Width in the formula when I did it thanks all !